# It’s complicated: The relationship of algorithmic fairness and the non‑discrimination provisions for high‑risk systems in the EU AI Act

### v5

@@ 3.1 The EU AI Act: History, Scope, and Definitions
Emergent technology can benefit from specific regulation. As one of the first comprehensive AI regulations worldwide (wodi2024artificial), the AI Act (AIA) introduces harmonised rules and provisions for the use and “placing on the market” of AI systems within the European Union (Article 1(2) AIA). The AI Act is a mix of a product safety regulation and fundamental rights protection (almada2023eu), aiming to establish a level playing field for AI technologies across the Union (nolte2024robustness). Notably, the original drafts of the European Commission did not include explicit provisions on individual rights (hacker2024generative). At that stage, individual rights — understood as protecting the rights of individuals affected — were largely absent. In fact, Members of the European Parliament initially did not prioritise the regulation of algorithmic discrimination (chiappetta2023navigating). However, over the course of the legislative process, different aspects related to fairness were integrated into the final text.
@@
The AI Act relies in part on the New Legislative Framework of the European Union (EU) (kaminski2025american). The New Legislative Framework is a cornerstone of modern product safety law in the EU (see also (noauthor_commission_2022) for further details), which also applies, for example, to medical devices and children’s toys [^8] See https://single-market-economy.ec.europa.eu/single-market/goods/new-legislative-framework_en for an overview..
@@
The AI Act regulates AI systems and also includes provisions for General Purpose AI models. The relationship or distinction between a model and a system in the context of the AIA remains legally unresolved. Article 3(1) defines an “AI system” as “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments”, but the article does not provide a definition of “AI model”. Recital 97 clarifies that “although AI models are essential components of AI systems, they do not constitute AI systems on their own. AI models require the addition of further components, such as for example a user interface, to become AI systems.” This formulation does not conclusively settle the terminological relationship between models and systems, as discussed in more detail in (nolte2024robustness). Moreover, the regulation of “AI systems” contrasts with the (trustworthy) computer science literature, which typically studies machine learning models. Since our focus is the legal framework, we follow the terminology of the AI Act, even though computer science literature tends to focus on models rather than systems.
@@
The core regulatory structure of the AI Act is built on a risk-based approach (hacker2024regulating). Risk is defined in Article 3(2) AIA as “the combination of the probability of the occurrence of harm and the severity of that harm.” Based on this definition, the AI Act categorises AI systems into different risk levels, each associated with a specific set of regulatory requirements [^9] This classification has been subject to criticism, see e.g., (bosoer2023non).:
@@ 3.3 High-risk AI non-discrimination provisions
What is a high-risk AI system? Article 6 AIA defines different aspects of high risk. For example, according to Article 6(2) and in combination with Annex III, high-risk systems are, for example, intended to be used as safety components in the management and operation of critical digital infrastructure (Annex III(2) AIA) or systems intended to be used for the recruitment or selection of natural persons (Annex III(4)(a) AIA). The reasoning behind the specific rules in Article 6 AIA is that systems which pose a specific risk to fundamental rights or safety need tighter regulation (fraser2024acceptable) than other, less risky systems. Recital 48 states that the adverse impact caused by the AI system on fundamental rights is of particular relevance when classifying an AI system as a high-risk system. It is estimated that 5%-15% [^10] For a critical assessment of these numbers see (almada2023eu). of all AI systems fall under the category of high-risk systems (noauthor_commission_2021).
@@
Article 9 AIA: Non-discrimination provisions within risk management systems Article 9 of the AI Act covers risk management systems of high-risk systems. According to Article 9(2)(a-d) AIA, risk management systems are designed to identify, evaluate, and mitigate risks of AI systems in a continuous, iterative process throughout the entire lifecycle of a high-risk AI system. All risks after mitigation, called residual risks, need to be judged acceptable (soler2023analysis).
Of particular relevance to non-discrimination law is Article 9(2)(a) AIA, which requires the identification and analysis of known and reasonably foreseeable risks [^12] See for the definition of a foreseeable risk also the EU Blue Guide of the New Legislative Framework noauthor_commission_2022. that the high-risk AI system can pose to health, safety, or fundamental rights. Through the link to fundamental rights, indirectly, Article 9 AIA already requires the identification and analysis of non-discrimination issues (zuiderveen2024non). The initial draft of the AI Act did not include the reference to fundamental rights, and it was added by the EU Council in 2022131313See Council of the European Union, “Proposal for a Regulation of the European Parliament and the Council laying down harmonised rules on artificial intelligence (Artificial Intelligence Act) and amending certain Union legislative acts – General approach” Doc. 14954/22, 25 November 2022., highlighting the legislator’s intent to include non-discrimination aspects in the AI Act.
@@
The first observation we make is that the input data of an AI system should undergo a bias analysis. Most importantly, the term bias is neither defined in the AI Act nor is there a common understanding of its meaning (van2024using). It seems that the regulators had a more technical definition of bias in mind, focusing on the diversity of training data in different dimensions compared to social, ethical, or structural biases (hacker2024generative). This could imply difficulties in determining the regulatory content, also for the later standardisation process.
@@
Furthermore, the wording of Article 10(2)(f) AIA could imply an expost view: the provider has to know whether a bias will likely lead to discrimination. However, especially for larger models, it is technically very challenging, if not impossible, to predict the individual output of AI models (hacker2024generative; black2022model).
Article 10(2)(g) AIA: Appropriate bias detection, prevention, and mitigation on input data must include factors beyond technical means. Article 10(2)(f) AIA is complemented by Article 10(2)(g) AIA. Article 10(2)(g) AIA demands that “appropriate measures to detect, prevent and mitigate possible biases identified according to point (f)” must be implemented. Thus, Article 10 AIA not only mandates the examination of biases that lead to discrimination but also the mitigation of these biases. The bias tests must be documented and disclosed according to Article 11 AIA, along with Annexes IV and IXa AIA (zuiderveen2024non). It has been argued that with this focus on the input side, the AI Act seeks to remedy the root cause of biases (zuiderveen2024non).
Article 10(2)(g) AIA appears to be inspired by the technical literature on fairness metrics for algorithmic outputs. Nevertheless, Article 10(2)(g) AIA targets the input of algorithms. Therefore, classical fairness metrics for algorithmic outputs do not apply directly. Methods for detecting, preventing, and mitigating input bias remain relatively underexplored in the existing literature, and these methods are not clearly defined in the AI Act (wachter2024limitations; deck2024mapping).
@@
Article 10(5) AIA: A legal basis for the processing of personal data in non-discrimination contexts. Finally, Article 10(5) AIA allows the processing of Article 9 GDPR data (see also Recital 70 AIA). Article 9 GDPR protects special categories of personal data, such as genetic, biometric, or health data (Article 9(1) GDPR). There is a tension (deck2024implications) between the need for debiasing AI algorithms and data protection law, which Article 10(5) AIA aims to address. In order to effectively mitigate biases in AI systems, the processing of personal data (for example, to compute fairness metrics) is important (van2024using). Notably, this exception only applies in the high-risk regime. Thus, this exception is not applicable to non-high-risk systems. Nevertheless, Article 10(5) is an example of how the AI Act not only imposes burdens on providers of AI systems but also grants rights to these groups. This aspect is frequently underappreciated in the discourse regarding the AI Act.
@@
Article 15 (4) AIA: Output bias-related provisions in cases of feedback loops. In contrast to Article 10(2)(f) AIA, Article 15(4) AIA mandates that “High-risk AI systems that continue to learn after being placed on the market or put into service shall be developed in such a way as to eliminate or reduce as far as possible the risk of possibly biased outputs influencing input for future operations (feedback loops), and as to ensure that any such feedback loops are duly addressed with appropriate mitigation measures.” In this case, the output of a system is regulated, in contrast to the input of an AI system in Article 10 AIA. The reasoning behind this is that AI systems should not become echo chambers for biases (zuiderveen2024non). According to Recital 67 AIA, this is a particular concern when examining historical biases. It is important to note that this output provision only applies if the systems continue to learn after being placed on the market or put into service.
Other articles regarding high-risk systems requirements. While the previously discussed articles provide at least some intuition about the addressed concerns, we categorise the remaining articles of Table 1. Most of these paragraphs only contain links to fundamental rights, which show a limited impact on non-discrimination requirements. Thus, the discussion here is rather short. Article 13(3)(b)(iii) AIA mandates that information on the foreseeable misuse and the impact on fundamental rights be provided. Article 14(2) AIA mandates that human oversight is needed to minimise risks to fundamental rights. Article 17(1)(i) AIA and Article 26(5) AIA are related to reporting and information on serious incidents. Serious incidents are linked to discrimination through Article 3(49)(c) AIA: “serious incident” means an incident or malfunctioning of an AI system that directly or indirectly leads to any of the following: […] the infringement of obligations under Union law intended to protect fundamental rights”. Finally, Article 27(1) AIA requires a fundamental rights impact assessment when the high-risk system is used in specific systems within bodies governed by public law. Yet again, significant gaps exist between theoretical and methodological elaboration of these impact assessments (mantelero2024fundamental). Article 27(1) AIA only addresses public parties. A fundamental rights impact assessment is not necessary for private parties.
@@
Summary: High-risk systems balancing between specificity and generality. The AI Act aims to be a specific law for AI (zHodi2024conflict). For the ML and Law community, it might be challenging to understand the implications of this specific AI law. Some articles mention non-discrimination directly, while others focus on fundamental rights. Only Article 10 AIA and Article 15 AIA explicitly mention non-discrimination and biases, compared to the statement of fundamental rights protection in Article 9 AIA. However, Article 10 AIA only considers the input phase, while Article 15 AIA only addresses the output phase for feedback loops. It is unclear why the AI Act does not consistently regulate direct, indirect, and intersectional discrimination across both the input and output phases.
Furthermore, it is essential to note that most provisions predominantly focus on the providers of AI systems from the perspective of a product safety regulation. However, providers may not always be aware of the specific use cases in which a high-risk system will be deployed (correa2025better). Consequently, fairness testing conducted solely by the provider, regardless of the system development phase, may be insufficiently comprehensive or contextually inappropriate.
@@ 4 The role of standardisation in the AI Act
According to the New Legislative Framework, all AI systems and models entering the market require a conformity assessment demonstrating that AI systems and models comply with the provisions of the AI Act (kaminski2025american). This process will culminate in the establishment of harmonised technical standards (ebers2021standardizing).
The details of the conformity assessments are set out in Articles 40 ff. AIA [^14] For the process of establishing a harmonised standard see (kaminski2025american). . Depending on the specific product, either a self-assessment or a third-party assessment is possible (Article 43 AIA). The AI Act stipulates the requirements for the conformity assessments themselves but does not prescribe the specific content of these assessments (kaminski2025american).
Under Articles 40 and following, mandates have been issued to the European standardisation organisations. In March 2023, the European Commission issued Mandate M/935 to CEN (European Committee for Standardisation) and CENELEC JTC 21 (European Committee for Electrotechnical Standardisation) (kilian2025european). Thus, the details of non-discrimination procedures — especially concerning methods and the entire risk management framework outlined in Article 17 — will be developed through the standardisation process. A similar approach exists for GPAI models. For GPAI models, providers could follow the developed Code of Practice to show compliance with the AI Act (Article 55(2) AIA). Throughout the standardisation, experts and stakeholders will collaboratively refine non-discrimination measures for high-risk systems.
However, although standardisation committees have considerable discretion, they must operate within the bounds of legal norms. For example, if the AI Act demands discrimination testing for input data (and excludes output data), standardisation cannot necessarily go beyond this, though it may at times do so. Thus, standardisation in the context of algorithmic fairness can also produce a false sense of safety (laux2024three).
@@ 5 Interplay between traditional EU non-discrimination law and the AI Act for high-risk systems
With the adoption of the AI Act, questions arise as to how its regime will interact with pre-existing non-discrimination law in the EU. Whereas the AI Act primarily regulates the input side for high-risk systems, much of the machine learning research community’s work on fairness concerns algorithmic outputs. The AI Act could be regarded as a lex specialis (craig2011eu) so it prevails over general frameworks to the extent that it covers an issue; where it does not address certain discriminatory outcomes, however, the more general non-discrimination law remains applicable. Since the AI Act does not directly regulate algorithmic outputs for standard high-risk systems, traditional EU non-discrimination law may still apply in these cases.
@@
The relationship between the AI Act and existing fundamental rights protections, including those concerning non-discrimination, requires further clarification and research (lewis2025mapping). This also includes an analysis of when fundamental rights are vertically and horizontally applicable. Furthermore, the fundamental rights in the AI Act need to be balanced (kusche2024possible), and the material scope is limited due to the enumeration of protected attributes (xenidis2020eu). As pointed out, “gaps in the regulatory framework have left fundamental rights inconsistently protected in the AI Act.” (palmiotto2025ai) Nevertheless, since the AI Act does not provide protection for individuals regarding the output of algorithms, there is room for the applicability of the CFR. It remains an open research question how the protection in the CFR and the AI Act relate (lewis2025mapping).
@@
Secondary EU law, such as Directives 2004/113/EC and 2000/43/EC, provides more concrete protections against discrimination, and these instruments have clear horizontal applicability, unlike the frequently debated scope of the CFR (xenidis2020eu; lewis2025mapping). For example, Directives 2004/113/EC (Equal Treatment in Goods and Services Directive) and 2000/43/EC (Race and Ethnicity Equality Directive) specifically target non-discrimination regulations. These regulations address specific cases in which discrimination can occur. Due to their specificity, they have a limited material and personal scope as a result of the enumeration of protected attributes (xenidis2020eu; circiumaru2024discrimination). The development of the relationship between existing regulations and the AI Act, and the stance that national and EU courts will ultimately adopt, is still unresolved. In the past, the ECJ has already issued important rulings on how the non-discrimination provisions in Article 21 CFR must be interpreted in relation to secondary EU law (xenidis2020eu). It is very likely that in the future, the ECJ will continue to issue rulings to clarify the interplay between the EU AI Act, fundamental rights, and other directives and regulations.
@@
The AI Act holds the potential to complement traditional non-discrimination legislation. Traditional non-discrimination law primarily targets the protection of individuals. The EU AI Act and its rules on high-risk systems are fundamentally aligned with product safety regulation rather than human rights protection. Despite the identified challenges associated with this product-safety approach of the AI Act, it nonetheless provides a degree of protection. However, such an emphasis on products can overlook broader regulatory dimensions. In the context of EU data protection, it has been posited that a dual approach involving both individual protection and collaborative governance is necessary (kaminski2018binary). A dual approach might also be beneficial for non-discrimination law in the context of AI.

### v6

@@ 3.1 The EU AI Act: History, Scope, and Definitions
Emergent technology can benefit from specific regulation. As one of the first comprehensive AI regulations worldwide (Wodi, 2024), the AI Act (AIA) introduces harmonised rules and provisions for the use and “placing on the market” of AI systems within the European Union (Article 1(2) AIA). The AI Act is a mix of a product safety regulation and fundamental rights protection (Almada and Petit, 2023), aiming to establish a level playing field for AI technologies across the Union (Nolte et al., 2024). Notably, the original drafts of the European Commission did not include explicit provisions on individual rights (Hacker et al., 2024b). At that stage, individual rights — understood as protecting the rights of individuals affected — were largely absent. In fact, Members of the European Parliament initially did not prioritise the regulation of algorithmic discrimination (Chiappetta, 2023). However, over the course of the legislative process, different aspects related to fairness were integrated into the final text.
@@
The AI Act relies in part on the New Legislative Framework of the European Union (EU) (Kaminski and Selbst, 2025). The New Legislative Framework is a cornerstone of modern product safety law in the EU (see also (Commission, 2022) for further details), which also applies, for example, to medical devices and children’s toys [^8] See https://single-market-economy.ec.europa.eu/single-market/goods/new-legislative-framework_en for an overview..
@@
The AI Act regulates AI systems and also includes provisions for General Purpose AI models. The relationship or distinction between a model and a system in the context of the AIA remains legally unresolved. Article 3(1) defines an “AI system” as “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments”, but the article does not provide a definition of “AI model”. Recital 97 clarifies that “although AI models are essential components of AI systems, they do not constitute AI systems on their own. AI models require the addition of further components, such as for example a user interface, to become AI systems.” This formulation does not conclusively settle the terminological relationship between models and systems, as discussed in more detail in (Nolte et al., 2024). Moreover, the regulation of “AI systems” contrasts with the (trustworthy) computer science literature, which typically studies machine learning models. Since our focus is the legal framework, we follow the terminology of the AI Act, even though computer science literature tends to focus on models rather than systems.
@@
The core regulatory structure of the AI Act is built on a risk-based approach (Hacker et al., 2024a). Risk is defined in Article 3(2) AIA as “the combination of the probability of the occurrence of harm and the severity of that harm.” Based on this definition, the AI Act categorises AI systems into different risk levels, each associated with a specific set of regulatory requirements [^9] This classification has been subject to criticism, see e.g., (Bosoer et al., 2023).:
@@ 3.3 High-risk AI non-discrimination provisions
What is a high-risk AI system? Article 6 AIA defines different aspects of high risk. For example, according to Article 6(2) and in combination with Annex III, high-risk systems are, for example, intended to be used as safety components in the management and operation of critical digital infrastructure (Annex III(2) AIA) or systems intended to be used for the recruitment or selection of natural persons (Annex III(4)(a) AIA). The reasoning behind the specific rules in Article 6 AIA is that systems which pose a specific risk to fundamental rights or safety need tighter regulation (Fraser and y Villarino, 2024) than other, less risky systems. Recital 48 states that the adverse impact caused by the AI system on fundamental rights is of particular relevance when classifying an AI system as a high-risk system. It is estimated that 5%-15% [^10] For a critical assessment of these numbers see (Almada and Petit, 2023) of all AI systems fall under the category of high-risk systems (Commission, 2021).
@@
Article 9 AIA: Non-discrimination provisions within risk management systems Article 9 of the AI Act covers risk management systems of high-risk systems. According to Article 9(2)(a-d) AIA, risk management systems are designed to identify, evaluate, and mitigate risks of AI systems in a continuous, iterative process throughout the entire lifecycle of a high-risk AI system. All risks after mitigation, called residual risks, need to be judged acceptable (Soler Garrido et al., 2023).
Of particular relevance to non-discrimination law is Article 9(2)(a) AIA, which requires the identification and analysis of known and reasonably foreseeable risks [^12] See for the definition of a foreseeable risk also the EU Blue Guide of the New Legislative Framework Commission (2022). that the high-risk AI system can pose to health, safety, or fundamental rights. Through the link to fundamental rights, indirectly, Article 9 AIA already requires the identification and analysis of non-discrimination issues (Borgesius et al., 2024). The initial draft of the AI Act did not include the reference to fundamental rights, and it was added by the EU Council in 2022131313See Council of the European Union, “Proposal for a Regulation of the European Parliament and the Council laying down harmonised rules on artificial intelligence (Artificial Intelligence Act) and amending certain Union legislative acts – General approach” Doc. 14954/22, 25 November 2022., highlighting the legislator’s intent to include non-discrimination aspects in the AI Act.
@@
The first observation we make is that the input data of an AI system should undergo a bias analysis. Most importantly, the term bias is neither defined in the AI Act nor is there a common understanding of its meaning (van Bekkum, 2024). It seems that the regulators had a more technical definition of bias in mind, focusing on the diversity of training data in different dimensions compared to social, ethical, or structural biases (Hacker et al., 2024b). This could imply difficulties in determining the regulatory content, also for the later standardisation process.
@@
Furthermore, the wording of Article 10(2)(f) AIA could imply an expost view: the provider has to know whether a bias will likely lead to discrimination. However, especially for larger models, it is technically very challenging, if not impossible, to predict the individual output of AI models (Hacker et al., 2024b; Black et al., 2022).
Article 10(2)(g) AIA: Appropriate bias detection, prevention, and mitigation on input data must include factors beyond technical means. Article 10(2)(f) AIA is complemented by Article 10(2)(g) AIA. Article 10(2)(g) AIA demands that “appropriate measures to detect, prevent and mitigate possible biases identified according to point (f)” must be implemented. Thus, Article 10 AIA not only mandates the examination of biases that lead to discrimination but also the mitigation of these biases. The bias tests must be documented and disclosed according to Article 11 AIA, along with Annexes IV and IXa AIA (Borgesius et al., 2024). It has been argued that with this focus on the input side, the AI Act seeks to remedy the root cause of biases (Borgesius et al., 2024).
Article 10(2)(g) AIA appears to be inspired by the technical literature on fairness metrics for algorithmic outputs. Nevertheless, Article 10(2)(g) AIA targets the input of algorithms. Therefore, classical fairness metrics for algorithmic outputs do not apply directly. Methods for detecting, preventing, and mitigating input bias remain relatively underexplored in the existing literature, and these methods are not clearly defined in the AI Act (Wachter, 2024; Deck et al., 2024b).
@@
Article 10(5) AIA: A legal basis for the processing of personal data in non-discrimination contexts. Finally, Article 10(5) AIA allows the processing of Article 9 GDPR data (see also Recital 70 AIA). Article 9 GDPR protects special categories of personal data, such as genetic, biometric, or health data (Article 9(1) GDPR). There is a tension (Deck et al., 2024a) between the need for debiasing AI algorithms and data protection law, which Article 10(5) AIA aims to address. In order to effectively mitigate biases in AI systems, the processing of personal data (for example, to compute fairness metrics) is important (van Bekkum, 2024). Notably, this exception only applies in the high-risk regime. Thus, this exception is not applicable to non-high-risk systems. Nevertheless, Article 10(5) is an example of how the AI Act not only imposes burdens on providers of AI systems but also grants rights to these groups. This aspect is frequently underappreciated in the discourse regarding the AI Act.
@@
Article 15 (4) AIA: Output bias-related provisions in cases of feedback loops. In contrast to Article 10(2)(f) AIA, Article 15(4) AIA mandates that “High-risk AI systems that continue to learn after being placed on the market or put into service shall be developed in such a way as to eliminate or reduce as far as possible the risk of possibly biased outputs influencing input for future operations (feedback loops), and as to ensure that any such feedback loops are duly addressed with appropriate mitigation measures.” In this case, the output of a system is regulated, in contrast to the input of an AI system in Article 10 AIA. The reasoning behind this is that AI systems should not become echo chambers for biases (Borgesius et al., 2024). According to Recital 67 AIA, this is a particular concern when examining historical biases. It is important to note that this output provision only applies if the systems continue to learn after being placed on the market or put into service.
Other articles regarding high-risk systems requirements. While the previously discussed articles provide at least some intuition about the addressed concerns, we categorise the remaining articles of Table 1. Most of these paragraphs only contain links to fundamental rights, which show a limited impact on non-discrimination requirements. Thus, the discussion here is rather short. Article 13(3)(b)(iii) AIA mandates that information on the foreseeable misuse and the impact on fundamental rights be provided. Article 14(2) AIA mandates that human oversight is needed to minimise risks to fundamental rights. Article 17(1)(i) AIA and Article 26(5) AIA are related to reporting and information on serious incidents. Serious incidents are linked to discrimination through Article 3(49)(c) AIA: “serious incident” means an incident or malfunctioning of an AI system that directly or indirectly leads to any of the following: […] the infringement of obligations under Union law intended to protect fundamental rights”. Finally, Article 27(1) AIA requires a fundamental rights impact assessment when the high-risk system is used in specific systems within bodies governed by public law. Yet again, significant gaps exist between theoretical and methodological elaboration of these impact assessments (Mantelero, 2024). Article 27(1) AIA only addresses public parties. A fundamental rights impact assessment is not necessary for private parties.
@@
Summary: High-risk systems balancing between specificity and generality. The AI Act aims to be a specific law for AI (Ződi, 2024). For the ML and Law community, it might be challenging to understand the implications of this specific AI law. Some articles mention non-discrimination directly, while others focus on fundamental rights. Only Article 10 AIA and Article 15 AIA explicitly mention non-discrimination and biases, compared to the statement of fundamental rights protection in Article 9 AIA. However, Article 10 AIA only considers the input phase, while Article 15 AIA only addresses the output phase for feedback loops. It is unclear why the AI Act does not consistently regulate direct, indirect, and intersectional discrimination across both the input and output phases.
Furthermore, it is essential to note that most provisions predominantly focus on the providers of AI systems from the perspective of a product safety regulation. However, providers may not always be aware of the specific use cases in which a high-risk system will be deployed (Corrêa et al., 2025). Consequently, fairness testing conducted solely by the provider, regardless of the system development phase, may be insufficiently comprehensive or contextually inappropriate.
@@ 4 The role of standardisation in the AI Act
According to the New Legislative Framework, all AI systems and models entering the market require a conformity assessment demonstrating that AI systems and models comply with the provisions of the AI Act (Kaminski and Selbst, 2025). This process will culminate in the establishment of harmonised technical standards (Ebers, 2021).
The details of the conformity assessments are set out in Articles 40 ff. AIA [^14] For the process of establishing a harmonised standard see (Kaminski and Selbst, 2025). . Depending on the specific product, either a self-assessment or a third-party assessment is possible (Article 43 AIA). The AI Act stipulates the requirements for the conformity assessments themselves but does not prescribe the specific content of these assessments (Kaminski and Selbst, 2025).
Under Articles 40 and following, mandates have been issued to the European standardisation organisations. In March 2023, the European Commission issued Mandate M/593 to CEN (European Committee for Standardisation) and CENELEC JTC 21 (European Committee for Electrotechnical Standardisation) (Kilian et al., 2025) (updated to Mandate M/613 in 2025). Thus, the details of non-discrimination procedures — especially concerning methods and the entire risk management framework outlined in Article 17 — will be developed through the standardisation process. A similar approach exists for GPAI models. For GPAI models, providers could follow the developed Code of Practice to show compliance with the AI Act (Article 55(2) AIA). Throughout the standardisation, experts and stakeholders will collaboratively refine non-discrimination measures for high-risk systems.
However, although standardisation committees have considerable discretion, they must operate within the bounds of legal norms. For example, if the AI Act demands discrimination testing for input data (and excludes output data), standardisation cannot necessarily go beyond this, though it may at times do so. Thus, standardisation in the context of algorithmic fairness can also produce a false sense of safety (Laux et al., 2024).
@@ 5 Interplay between traditional EU non-discrimination law and the AI Act for high-risk systems
With the adoption of the AI Act, questions arise as to how its regime will interact with pre-existing non-discrimination law in the EU. Whereas the AI Act primarily regulates the input side for high-risk systems, much of the machine learning research community’s work on fairness concerns algorithmic outputs. The AI Act could be regarded as a lex specialis (Craig and De Búrca, 2011) so it prevails over general frameworks to the extent that it covers an issue; where it does not address certain discriminatory outcomes, however, the more general non-discrimination law remains applicable. Since the AI Act does not directly regulate algorithmic outputs for standard high-risk systems, traditional EU non-discrimination law may still apply in these cases.
@@
The relationship between the AI Act and existing fundamental rights protections, including those concerning non-discrimination, requires further clarification and research (Lewis et al., 2025). This also includes an analysis of when fundamental rights are vertically and horizontally applicable. Furthermore, the fundamental rights in the AI Act need to be balanced (Kusche, 2024), and the material scope is limited due to the enumeration of protected attributes (Xenidis and Senden, 2020). As pointed out, “gaps in the regulatory framework have left fundamental rights inconsistently protected in the AI Act.” (Palmiotto, 2025) Nevertheless, since the AI Act does not provide protection for individuals regarding the output of algorithms, there is room for the applicability of the CFR. It remains an open research question how the protection in the CFR and the AI Act relate (Lewis et al., 2025).
@@
Secondary EU law, such as Directives 2004/113/EC and 2000/43/EC, provides more concrete protections against discrimination, and these instruments have clear horizontal applicability, unlike the frequently debated scope of the CFR (Xenidis and Senden, 2020; Lewis et al., 2025). For example, Directives 2004/113/EC (Equal Treatment in Goods and Services Directive) and 2000/43/EC (Race and Ethnicity Equality Directive) specifically target non-discrimination regulations. These regulations address specific cases in which discrimination can occur. Due to their specificity, they have a limited material and personal scope as a result of the enumeration of protected attributes (Xenidis and Senden, 2020; Cîrciumaru, 2024). The development of the relationship between existing regulations and the AI Act, and the stance that national and EU courts will ultimately adopt, is still unresolved. In the past, the ECJ has already issued important rulings on how the non-discrimination provisions in Article 21 CFR must be interpreted in relation to secondary EU law (Xenidis and Senden, 2020). It is very likely that in the future, the ECJ will continue to issue rulings to clarify the interplay between the EU AI Act, fundamental rights, and other directives and regulations.
@@
The AI Act holds the potential to complement traditional non-discrimination legislation. Traditional non-discrimination law primarily targets the protection of individuals. The EU AI Act and its rules on high-risk systems are fundamentally aligned with product safety regulation rather than human rights protection. Despite the identified challenges associated with this product-safety approach of the AI Act, it nonetheless provides a degree of protection. However, such an emphasis on products can overlook broader regulatory dimensions. In the context of EU data protection, it has been posited that a dual approach involving both individual protection and collaborative governance is necessary (Kaminski, 2018). A dual approach might also be beneficial for non-discrimination law in the context of AI.